OpenStudy (anonymous):

Does the infinite geometric series diverge or converge? 1/6+1/12+1/24+1/48+... a.It diverges;it does not have a sum b.It diverges;it has a sum c.It converges;it has a sum d. It converges;it does not have a sum

OpenStudy (anonymous):

\[\frac{ 1 }{ 6 }+\frac{ 1 }{ 12 }+\frac{ 1 }{ 24 }+\frac{ 1 }{ 48 }+....\]

OpenStudy (anonymous):

This geometric series has ratio 1/2 DEFINITION:If r is between −1 and +1, the terms of the series become smaller and smaller, approaching zero in the limit and the series converges to a sum

OpenStudy (anonymous):

So the correct answer is "C."